OpenStudy (anonymous):

Shannon makes large cookies. These cookies are 22 inches in diameter. How much icing will she need to cover the top of each cookie?

OpenStudy (rulnick):

enough to cover pi * 22^2 / 4 in^2, or about 380.1 sq in. of area.

OpenStudy (anonymous):

Area of cookie = πR^2 22 22 * 121 = ----- * 11^2 = --------- = 380.29 SQ.IN. 7 7